OpenStudy (anonymous):

] Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1

OpenStudy (campbell_st):

let y = 3 - 2x then substituting into the 2nd equation -3x + 4( 3 - 2x) = 1 or -3x + 12 - 8x = 1 -11x + 12 = 1 solve for x.... when you have a value for x, substitute it into the 1st equation to find y this will be the ordered pair solution